1) Fill in the blanks: a. The principal quantum number,n, can have integer values from b. The angular momentum quantum #,
Answer #1

voo as we Date: Page (a 3 d 5 f7 1 to x . Explain = l = 0 - 5 e = 1 - P o to n-1 l= 2 - d l=3 - F l to tl value of l = n-1 Ex

(e 1 cone) 10 (-5,- 4, -3,- 2, 1, 0, +1, +2, +3, +4, +5 (10 value (nx27 totul value of el n2 = orbital Number) 16 (h) Magneti

thanks